OpenStudy (anonymous):

how do you find the intercept of the equation 19x + y = 0

OpenStudy (auctoratrox):

you rewrite the equation in slope-intercept form, y=mx+b. in this case, it is y = -19x. you can now input y=o to get the x-intercepts and x=0 to get the y-intercepts.

OpenStudy (anonymous):

okay thanks, oh and how would you find the slope of this equation?

OpenStudy (auctoratrox):

m is the slope since we're using slope-intercept form

OpenStudy (anonymous):

so in this case the slope would be -19?

jimthompson5910 (jim_thompson5910):

you are correct, the slope is -19

OpenStudy (anonymous):

ok cool..thanks

OpenStudy (anonymous):

okay how would you determine the angle of intersection for the lines: 4x-y+2=0 and 19x+y=0

OpenStudy (anonymous):

i know you would have to find the slope of both equations which i did but i'm getting \[\theta=17 when \it should be \theta=163\]

jimthompson5910 (jim_thompson5910):

notice how 17+163 = 180, so you're just calculating the supplement to the actual angle that you want

OpenStudy (anonymous):

ohhh thats what i didn't do..thanks once again

jimthompson5910 (jim_thompson5910):

you're welcome